# Cascaded Self-Evaluation Augmented Training for Efficient Multimodal Large Language Models

### v1

@@ 5 Conclusion
Limitation
Potential limitations may include a small amount of resource consumption. The method is designed specifically for EMLLMs, and it may not currently be the optimal solution for larger MLLMs.
@@ Appendix A Appendix
A.1 Supplementary Experiments
A.1.1 Case Study
Figure 6: A table analysis sample in MathVista. The meanings of the colors are the same as those in Figure 5.
Figure 6 is a supplementary case, the observed findings can be referred to the section 4.2.3.
A.1.2 Datasets
@@
Table 5: Statistics of Datasets.
Dataset MathV360k CoT Cas-SEAT
#Samples 339k 160k 167k
A.1.3 Hyperparameters and Training Schedules
@@
Table 6: Hyperparameters and training schedules of LLaVA.
@@
Table 7: Hyperparameters and training schedules of Qwen2-VL.

### v2

@@ Appendix A Appendix
A.1. Supplementary Experiments
A.1.1. Cost Analysis
As shown in Table 9, DDF significantly reduces the training data size and consequently the training duration while using the same LLaVA-v1.5-7B model. Our method makes LLaVA-v1.5-7B far outperform LLaVA-v1.5-13B. In this context, we further compared the inference cost of LLaVA-v1.5-7B under our method with that of LLaVA-v1.5-13B, demonstrating the efficiency of our approach.
A.1.2. Datasets
@@
A.1.3. Hyperparameters and Training Schedules
@@
Table 7. Hyperparameters of LLaVA-1.5.
@@
Table 8. Hyperparameters of Qwen2-VL.
@@
Table 9. Comparison of training and inference cost.
Training speed
On raw dataset 18h
On Cas-SEAT-DDF dataset 10h
Inference speed
LLaVA-v1.5-7B Input: 7700 tokens/s Output: 100 tokens/s
LLaVA-v1.5-13B Input: 3900 tokens/s Output: 50 tokens/s
Table 10. Statistics of Datasets.
Dataset MathV360k CoT Cas-SEAT
#Samples 339k 160k 167k
Table 11. Comparison of Cas-SEAT and the Baselines on Math-V Dataset. ALG, ARI, CG, and COM respectively denote Algebra, Arithmetic, Combinatorial Geometry, and Combinatorics.
Math-V
Model Extra Data All Level1 Level2 Level3 Level4 Level5 ALG ARI CG COM
Heuristics Baselines
Random Choice Base 0.0717 - - - - - 0.0150 0.0710 0.0970 0.0480
Close-source Models Inference
Qwen-VL-Plus Base 0.1072 - - - - - 0.1130 0.1430 0.1270 0.0480
Qwen-VL-Max Base 0.1559 - - - - - 0.1070 0.2000 0.1690 0.1250
GeminiPro Base 0.1766 - - - - - 0.1510 0.2070 0.2010 0.1190
GPT4V Base 0.2276 - - - - - 0.2730 0.3570 0.2110 0.1670
Open-Source Models Inference
SPHINX(V2) Base 0.0970 - - - - - 0.0670 0.1290 0.0750 0.0770
ShareGPT4V-7B Base 0.1053 - - - - - 0.0550 0.1290 0.1010 0.0480
LLaVA-v1.5-13B Base 0.1112 - - - - -
[…]
Finetune 0.1743 0.2075 0.1951 0.0893 0.1778 0.1912 0.1053 0.0000 0.2632 0.0000
CoT 0.1414 0.2075 0.1951 0.0893 0.1778 0.1912 0.0526 0.0000 0.3684 0.0526
SEAT 0.0757 0.1400 0.0690 0.0364 0.0444 0.0896 0.1053 0.0000 0.2632 0.0000
[…]
Open-Source Models Self-Evaluation
Base 0.0757 0.1509 0.1098 0.0714 0.2222 0.1765 0.0000 0.0000 0.3529 0.1053
Finetune 0.1776 0.2075 0.1951 0.0893 0.1778 0.2059 0.1053 0.0000 0.2632 0.0000
CoT 0.1447 0.1321 0.1341 0.2321 0.1778 0.1912 0.1053 0.0000 0.3684 0.0526
[…]
LLaVA-v1.5-7B Cas-SEAT 0.2763 0.2642 0.2439 0.2500 0.3111 0.3235 0.3158 0.1579 0.4211 0.2105
Improve 55.57% 27.33% 25.01% 7.71% 40.01% 57.12% 100.00% 0.00% 14.31% 33.31%
A.1.4. Evaluation on Different EMLLMs
Figure 5. Comparison of Cas-SEAT and the Baseline based on Qwen2-VL(2B).
A.1.5. Case Study
Figure 6 is a supplementary case, the observed findings can be referred to the section 4.2.2. Figure 7 is a reject sample case, which is detrimental to the training of EMLLMs.
Figure 6. A table analysis sample in MathVista. The meanings of the colors are the same as those in Figure 4.
Figure 7. A case of the reject samples.